Suppose that the number of asbestos particles in a sample of 1 squared centimeter of dust is a Poisson random variable with a mean of 1000.

What is the probability that 10 squared centimeters of dust contains more than 10120 particles? Use normal approximation without continuity correction. Round your answer to 3 decimal places.
Answer #1

solution;

Given data:

as mean =1000 and std deviation =(1000)1/2

also for 10 square cm . mean =1000*10 =10000

and std deviation =(1000)1/2 *(10)1/2 =100

hence P(X>10120) =P(Z>(10120-10000)/100) =P(Z>1.2) =1-P(Z<1.2) =1-0.8849 =0.1151

please give me thumb up